blog:20241222

20241222

1618-0337

Dimitry, Sarah, Saisri, Roy, Percy, Bhanu

Dome, 433

Variable stars URA

Scope Camera Focus
MainASI 26002065.1
FinderASI 482?
Time Observed Source Name Filter Int Time [s] Gain Airmass Mag (filter) RA [h] DEC [deg] Source Type Scope Purpose Notes
20241222T1924 2MASS_J00263594+6047535 V 90s 150 airmass mag(V) 00 60 Variable main URA tracking? Others?
20241222T2002 HD_236948 V 90s 150 airmass mag(V) 02 58 Variable main URA tracking? Others?
20241222T2028 2MASS_J00263594+6047535 V 90s 150 airmass mag(V) 00 68 Variable main URA tracking? Others?
20241222T2051 HD_236948 V 120s 150 airmass mag(V) 02 58 Variable main URA tracking? Others?
20241222T2115 HD_236948 V 90s 150 airmass mag(V) 02 58 Variable main URA tracking? Others?
20241222T2137 2MASS_J00263594+6047535 V 60s 150 1.236 mag(V) 00 60 Variable main URA tracking? Others?
20241222T2155 HIP_10332 V 120s 150 1.103 mag(V) 02 58 Variable main URA tracking? Others?
20241222T2210 2MASS_J02343129+5849537 V 60s 150 1.180 mag(V) 02 58 Variable main URA tracking? Others?
20241222T2334 HIP_11420 V 60s 150 1.197 mag(V) HH DD Variable main URA tracking? Others?
20241222T2341 HD_236948 V 60s 150 1.261 mag(V) HH DD Variable main URA tracking? Others?
20241222T0014 2MASS_J00263594+6047535 V 90s 150 1.678 mag(V) 00 60 Variable main URA tracking? Others?
20241222T0040 HIP_10332 V 120s 150 1.454 mag(V) 02 58 Variable main URA tracking? Others?
20241222T0059 HD_236948 V 90s 150 1.464 mag(V) 02 58 Variable main URA tracking? Others?
20241222T0120 HIP_11420 V 60s 150 1.481 mag(V) HH DD Variable main URA tracking? Others?
20241222T0138 2MASS_J02343129+5849537 V 60s 150 1.528 mag(V) 02 58 Variable main URA tracking? Others?
20241222T0158 FLATS N/A 30s 150 N/A mag(V) HH DD Flats main Calibration tracking? Others?
20241222T0228 DARKS N/A 120s 150 N/A mag(V) HH DD Flats main Calibration tracking? Others?
20241222T0307 DARKS N/A 60s 150 N/A mag(V) HH DD Flats main Calibration tracking? Others?
20241222T0333 DARKS N/A 90s 150 N/A mag(V) HH DD Flats main Calibration tracking? Others?

Evening Shift

  1. Mirror doors
    1. Only south door opened
    2. –> Closed door, tried again, all doors opened :)
  2. DFMTCS System Time was off by two minutes
    1. Dimitry updated system time to match cellphone
  3. Dome was misaligned to telescope
    1. Did not fix immediately (forgot)
  4. Pointed at Deneb
    1. Was off, so used SharpCap plate solving to find RA/DEC (J2000.0)
    2. Updated DFM TCS (Telescope > Initialization)
    3. Success
  5. Moved to HD 201935 for focusing
    1. Noticed immediately that tracking was off
    2. Took 10 frames 5s apart and stored on desktop
    3. Roy will analyze tonight to poke at tracking
    4. In the meantime, used PHD2 to guide
  6. Moved to 2MASS_J00331619+6254260
    1. Could not identify variable/reference star in finder chart
    2. Tried moving around, a range of camera settings, comparing to SIMBAD. Gave up after 20 minutes.
  7. Moved to 2MASS_J00263594+6047535 (aiming for a brighter variable)

Midnight Shift

  • 2100 - First capture of 2MASS_J02343129+5849537
    • Tracking switch was in auxiliary position, set it back and tracking is good now
  • 2234 - First capture of HIP_11420
  • 2341 - Third capture of HD_236948
    • Realized we used the wrong exposure time, will retake
  • 0014 - Fourth capture of 2MASS_J00263594+6047535
    • 90s exp time
    • Nearly saturating at this point
    • No more captures of this target
  • 0040 - Third capture of HIP_10332
    • 120s exp time
  • 0059 - Retaking 3rd capture of HD_236948
    • Correct exp time - 90s
  • 0120 - Second capture of HIP_11420
    • 60s exp time
  • 0138 - Second set of frames for 2MASS_J02343129+5849537
    • 60s exp time
  • 0156 - 30s flats
  • 0202 - Accidental 2nd set of 30s flats
  • 0202 - Shutting down the dome for darks
  • 0213 - North and south mirror doors not closing
  • 0238 - both mirror doors closed
    • broke thermalization, letting heat into the dome and thawing the mirror door motors enough for the south door to close on its own
    • Bhanu hand-spun the north mirror door closed
  • 0257 - First set of darks 120s
  • 0307 - Second set of darks, 60s
  • accidentally put the data computer into sleep mode, restarted
  • 0333 - Last set of darks (90s) done
  • Completed control room shutdown
  • Noticed exposure time discrepancy in the third set of captures for 2MASS_J00263594+6047535. Exp time is 60 instead of 90. Will use 90s exp time for the rest of the session
  • Percy needs to identify a clear naming convention for dark and flat frames and include in training
  • Percy accidentally took another set of 30s flats under the name darks_20241222_120s, please delete that one. Thank you!
  • In freezing temperatures, it may be necessary to break thermalization to close the mirror doors for flats. An alternative would be to keep the mirror doors open for darks and maintain thermalization until afterwards.
  • Add mirror door malfunctions to troubleshooting training
    • manually closing by hand-spinning motors
    • pointing the telescope to increase gravitational torque

cold, mostly clear, mostly.. 17 deg Fahrenheit - below freezing